### v2

# Preserve or Modify? Context-Aware Evaluation for Balancing Preservation and Modification in Text-Guided Image Editing

Abstract
The development of vision-language and generative models has significantly advanced text-guided image editing, which seeks the preservation of core elements in the source image while implementing modifications based on the target text. However, existing metrics have a context-blindness problem, indiscriminately applying the same evaluation criteria on completely different pairs of source image and target text, biasing towards either modification or preservation. Directional CLIP similarity, the only metric that considers both source image and target text, is also biased towards modification aspects and attends to irrelevant editing regions of the image. We propose AugCLIP, a context-aware metric that adaptively coordinates preservation and modification aspects, depending on the specific context of a given source image and target text. This is done by deriving the CLIP representation of an ideally edited image, that preserves the source image with necessary modifications to align with target text. More specifically, using a multi-modal large language model, AugCLIP augments the textual descriptions of the source and target, then calculates a modification vector through a hyperplane that separates source and target attributes in CLIP space. Extensive experiments on five benchmark datasets, encompassing a diverse range of editing scenarios, show that AugCLIP aligns remarkably well with human evaluation standards, outperforming existing metrics. The code will be open-sourced for community use.
\$*\$\$*\$footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
1 Introduction
Building on advancements in vision-language models (Radford et al., 2021; Li et al., 2022; Geng et al., 2023), recent generative models (Kawar et al., 2022; Brooks et al., 2022; Hertz et al., 2022) have been widely utilized as creative tools for image editing via text instructions. Text-guided image editing models enable the modification of images in response to textual guidance, ensuring that changes are aligned with the provided instructions. The primary objective of these models is to apply the necessary modifications guided by the target text while preserving most of the source image.
Figure 1: Context Blindness Problem of Existing Evaluation Metrics. Evaluation metrics should consider the specific context of a given source image and target text. However, existing metrics exhibit context-blindness, applying the same criteria of either ‘preserve’ (P) or ‘modify’ (M) across the entire image. Our proposed metric, AugCLIP, is a context-aware metric that flexibly applies different criteria to local regions of the image.
Despite the remarkable advancements in editing models, there has been a lack of rigorous evaluation methods, tailored specifically for text-guided image editing. Consequently, most studies (Hertz et al., 2023; Basu et al., 2023; Gal et al., 2022; Kim & Ye, 2021; Brooks et al., 2022; Ruiz et al., 2023; Kocasari et al., 2022) have heavily relied on human evaluation, which provides balanced consideration of preservation and modification aspects. However, as it is costly and impractical for real-world applications, researchers have adapted automatic evaluation metrics (Heusel et al., 2017; Zhang et al., 2018; Caron et al., 2021; Hessel et al., 2021) originally designed for other vision tasks, such as image generation or captioning.
In text-guided editing scenarios, it is essential to assess which elements in the source to preserve and which in the target to modify into, based on the specific context of the given source image and target text (see Fig. 1(a)-(b) for example). Existing metrics, however, exhibit context blindness, meaning preservation-centric metrics completely ignore the target context while modification-centric metrics do not account for the source context. As shown in Fig. 1(c), preservation- and modification-centric metrics apply a fixed standard across the entire image, regardless of the true editing requirement of the given context. Even combining them fails to overcome the context blindness problem.
Among existing metrics, directional CLIP similarity (Gal et al., 2022) is unique because it provides a method to consider both the source image and target text in the evaluation. Even this metric suffers from a context-blindness problem, using a fixed standard without considering the specific relationship between the source image and target text. We show that directional CLIP similarity tends to prioritize modification over preservation, and often focuses on peripheral parts rather than regions that are edited as guided by target text, in Sec. 3. These observations underscore the need for a context-aware metric to dynamically adjust the evaluation standard, balancing both aspects in response to diverse editing contexts.
Based on our comprehensive analysis, we propose a novel context-aware metric, AugCLIP, which evaluates the quality of the edited image by comparing it with an estimated representation of a well-edited image. To balance preservation and modification, a well-edited image is represented by a minimum modification on the source image, while matching the target text. We use multi-modal large language models (MLLMs) to extract attributes that capture various visual aspects of the source image and target text. With these attributes, we derive a hyperplane that separates the source and target to derive the ideal modification as a projection to this hyperplane. AugCLIP evaluates how closely the edited image aligns with the estimated ideal image in CLIP space.
Our metric AugCLIP demonstrates remarkable improvement in alignment with human judgment on diverse editing scenarios such as object, attribute, and style alteration compared to all other existing metrics. Moreover, our metric is even applicable to personalized generation, the DreamBooth dataset, where the objective is to identify an object in the source image and edit it into a completely novel context. This shows the flexibility of AugCLIP, which seamlessly applies to a variety of editing directions. Notably, our metric excels in identifying minor differences between the source image and the edited image, showing superb ability in complex image editing scenarios such as MagicBrush.
The major contributions are summarized as follows.
  • We point out the unreliability of current evaluation metrics in text-guided image editing, noting their inability to balance preservation and modification effectively due to inherent context blindness.
  • We introduce AugCLIP, a context-aware metric for image editing by automatically augmenting descriptions via MLLM and estimating a balanced representation of preservation and modification, which takes into account the relative importance of each description.
  • AugCLIP demonstrates a significantly high correlation with human evaluations across various editing scenarios, even in complex applications where existing metrics struggle.
2 Background on Existing Metrics
In the context of text-guided image editing, evaluation metrics should assess the quality of edited images in terms of both preservation of the source image and modification towards the target text.
Preservation-Centric Metrics
FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and L2 distance measure feature-level distance between images. These metrics are primarily used in image generation tasks to compare the distributions of real and generated images. When applied to editing scenarios, they assess the distance between the source and edited images, completely ignoring the target text guiding the editing process. Segmentation consistency (SC) (Kim & Ye, 2021) measures the structural change. Additionally, DINO similarity (DINO) and CLIP-I evaluate the semantic similarity between the source and edited images within the DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) embedding spaces, also focusing solely on preservation.
Modification-Centric Metrics
To assess how well an edited image aligns with the target text, the primary approach involves using a pre-trained multi-modal model (Radford et al., 2021). CLIPScore (CLIP-T) (Hessel et al., 2021) is the most widely adopted metric in this category, measuring the similarity between the edited image and the target text. This metric completely ignores the source image that should be considered for the preservation aspect.
Holistic Evaluation of Preservation and Modification
Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$) (Gal et al., 2022) is unique among existing metrics as it assesses both preservation and modification aspects by evaluating directional alignment between the original and edited images concerning the source and target text.
However, these metrics employ a fixed evaluation standard, disregarding the context of each editing scenario, which leads to context blindness. As discussed in Sec. 3, combining preservation- and modification-centric metrics or using directional CLIP similarity fails to balance preservation and modification adequately, resulting in misalignment with human judgment.
Metrics Used in Text-to-Image Generation
While text-to-image generation tasks benefit from a variety of evaluation metrics utilizing pre-trained multi-modal large language models (Hu et al., 2023; Li et al., 2024b; Lu et al., 2024; Huang et al., 2023; Cho et al., 2024) or VQA models (Hu et al., 2023; Ghosh et al., 2024), these metrics focus solely on prompt faithfulness. In contrast, image editing scenarios require a balance of image preservation and prompt adherence, making these metrics irrelevant to the evaluation of editing models.
3 Problems of Existing Metrics in Balancing Preservation and Modification Aspects
Figure 2: Combination of Preservation- and Modification-Centric Metrics Deteriorates in Performance. The plot shows the human alignment score ${\bm{s}}_{\text{align}}$ measured by a linear interpolation of {DINO, SC, CLIP-I} and CLIP-T. The results show that combining rather degrades the alignment with human judgment.
Existing evaluation metrics, that focus on either one of the preservation or modification aspects, face a challenge in assessing both aspects. Does combining these two metrics successfully balance the preservation and modification aspects? We systematically analyze how these metrics and their combinations still fail in balancing the two colliding objectives, suffering from context-blindness (Sec. 3.1). Moreover, we focus on directional CLIP similarity, the unique metric that accounts for both source image and target text, to reveal that it presents two significant problems that question its reliability (Sec. 3.2).
3.1 Combination of Preservation- and Modification-Centric Metrics
We explore whether combining preservation- and modification-centric metrics can improve alignment with human judgments across two benchmark datasets: EditVal (Basu et al., 2023) and CelebA (Liu et al., 2015) (Results on the other three benchmark datasets are deferred to the Appendix). To do so, we combine CLIP-T, a modification-centric metric, and preservation-centric metrics (LPIPS, DINO similarity, CLIP-I) with varying interpolation values $\gamma\in[0,1]$. Note that both metrics are scaled to the same range before interpolation, ensuring the interpolation value $\gamma$ is properly reflected in the final combined score.
As illustrated in Fig. 2, the combination of two powerful metrics in preservation and modification often leads to performance degradation. This manifests that a simple combination of the colliding metrics cannot be a reliable evaluation metric. This is due to the context-blindness of these metrics, failing to holistically consider source image and target text. Preservation-centric metrics compare only the source and edited images, ignoring the target context. Similarly, modification-centric metrics consider only the target text, overlooking the source context. This highlights the need for a tailored metric that is contextually aware of both preservation and modification, ensuring a reliable alignment with human evaluations across diverse datasets and tasks.
[…]
3.2 Directional CLIP Similarity
Directional CLIP similarity ($\mathrm{CLIP}_{\text{dir}}$) is designed to assess both preservation and modification in text-guided image editing by measuring the similarity between image and text directions as follows:
$\text{CLIP}_{\text{dir}}=\texttt{cs}(\Delta I,\Delta T)$ (1)
$=\texttt{cs}\Big{(}E(I_{\text{edit}})-E(I_{\text{src}}),E(T_{\text{trg}})-E(T_{\text{src}})\Big{)},$ (2)
where $\texttt{cs}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ denotes cosine similarity and $E(\cdot)$ is a CLIP encoder for either image or text. Here, $I_{\text{edit}}$ is the edited image to be evaluated, which is edited from the source image $I_{\text{src}}$ according to the target text $T_{\text{trg}}$. For evaluation of directional CLIP similarity, $I_{\text{src}}$, the corresponding text about the source image, is required. We analyze that $\mathrm{CLIP}_{\text{dir}}$ has two major problems that hinder the balanced evaluation of preservation and modification.
$\mathrm{CLIP}_{\text{dir}}$ overly emphasizes modification.
Directional CLIP similarity overly emphasizes the alignment with the target text while neglecting the preservation of the source image. Despite attempting to incorporate preservation by subtracting $T_{\text{src}}$ from $T_{\text{trg}}$, $\mathrm{CLIP}_{\text{dir}}$ frequently favors excessively modified images over ground truth ideally edited images as shown in samples in Fig. 3(a). Specifically, $\mathrm{CLIP}_{\text{dir}}$ metric fails on $63\%$ of the 100 editing cases in TEdBench (Kawar et al., 2022) and $35.1\%$ of the 1053 cases in MagicBrush (Zhang et al., 2024) (Details in Sec. 5). This problem happens since $\mathrm{CLIP}_{\text{dir}}$ is designed to assume that a well-edited image should primarily adhere to the target text. Thus, it blindly favors the visual features of the target text, leading to unreliable results.
$\mathrm{CLIP}_{\text{dir}}$ overlooks edited regions.
$\mathrm{CLIP}_{\text{dir}}$ often focuses on peripheral or unmodified local areas of the edited image, overlooking the specific areas that are edited by the text. Using the relevancy map (Chefer et al., 2021), ${\bm{R}}(\Delta I;\Delta T)$ [^1] The original relevancy map ${\bm{R}}$ utilizes an image-text pair. To adapt it to $\mathrm{CLIP}_{\text{dir}}$, we subtract two relevancy maps. Details are in the Appendix., to visualize attention on image regions corresponding to the text direction, we observe that $\mathrm{CLIP}_{\text{dir}}$ frequently fails to concentrate on the image regions that are directly relevant to the target text as shown in Fig. 3(b). The figure also shows that adding explicit visual details guides $\mathrm{CLIP}_{\text{dir}}$ to attend to edited regions properly.
Figure 4: (a) Description Extraction Process. The source image describes a young child standing on the balance board. Target text guides the editing model to make the girl sit. The source and target attributes are extracted with MLLM. (b) Evaluation Process of AugCLIP. The two edited images demonstrate i) An older woman sitting down with legs crossed ($I_{\text{edit}}^{1}$) and ii) A young girl sitting on the floor ($I_{\text{edit}}^{2}$). AugCLIP derives the ideal image representation as a minimum modification $\mathbf{v}$ on the source image to be classified as target. The second image that is closer to $I_{\text{src}}+\mathbf{v}$ shows a higher score, while the first image that is excessively modified with lost source identity shows a lower score.
4 AugCLIP: A Context-Aware Evaluation Metric
We introduce a novel evaluation metric, AugCLIP, for scoring the editing quality of the image in terms of preservation and modification. Given a target text $T_{\text{trg}}$ and a source image $I_{\text{src}}$, AugCLIP measures the cosine similarity between the edited image $I_{\text{edit}}$ and the ideal editing. Since the ideal edited image is not given in most evaluation cases, we estimate the representation of an ideally edited image as a modification on the source image, $I_{\text{src}}+\mathbf{v}$, where $\mathbf{v}$ is the modification vector. Then, the editing quality of the edited image $I_{\text{edit}}$ is evaluated as
$\texttt{AugCLIP}\left(I_{\text{edit}};I_{\text{src}},T_{\text{trg}}\right)=\texttt{cs}\left(E(I_{\text{edit}}),E(I_{\text{src}})+\mathbf{v}\right),$ (3)
where $E(\cdot)$ is a CLIP encoder.
The evaluation score ranges between $[0,1]$, where the maximum value $1$ indicates that the edited image is ideally modified by an editing model. We outline the framework of AugCLIP in Fig. 4. We extract the visual features of the source image and target text (Sec. 4.1), then derive the CLIP representation of an ideal modification that satisfies the two criteria: ‘i) Does the edited image follow the target text?’ and ‘ii) Does the edited image preserve the source image?’ (Sec. 4.2).
4.1 Extracting and Processing Attributes
Source & Target Attribute Extraction
As described in Fig. 4(a), we extract visual descriptions about attributes of the source image and target text using a state-of-the-art multi-modal large language model (MLLM), GPT-4V (OpenAI, 2023). From the source image, we extract visual features that can be the object of preservation. Likewise, from the target text, we extract visual features that can be the object of modification. The extraction process is instructed by five annotated in-context examples, which are provided in the Appendix.
Encoding into CLIP Space
After attribute extraction from MLLM, each attribute is encoded into CLIP space. The source attributes are denoted as $\mathcal{D}_{S}$ and target attributes as $\mathcal{D}_{T}$.
Relative Weighting on Attributes
In editing scenarios, some of the attributes are more important than others. For instance, drawing stripes is more essential than a short mane to represent a zebra. To prioritize this, we emphasize the similarity between key attributes within the same class. Since multi-modal large-language models sample based on the empirical probability, frequently extracted attributes indicate higher importance in representing the context. Formally, the importance of each visual attribute in $\mathcal{D}_{S}$ and $\mathcal{D}_{T}$ is represented as a cosine similarity between CLIP embeddings in the same set as
$\textit{Importance}(v_{i})=\sum_{j\neq i}\frac{u_{i}\cdot v_{j}}{\|u_{i}\|\|v_{j}\|},$ (4)
for all $u_{j}$ in the same attribute set, either $\mathcal{D}_{S}$ or $\mathcal{D}_{T}$.
Additionally, the importance should be down-weighted by how much the visual attribute hinders the preservation of the source image or modification of the target text. More specifically, a collision between source and target attributes is measured by averaged cosine distance to CLIP embeddings in the opposite set as
$\textit{Collision}(v_{i})=-\sum_{l}\frac{v_{i}\cdot w_{l}}{\|v_{i}\|\|w_{l}\|},$ (5)
for all $w_{l}$ in the opposite attribute set.
Adding these two measures, the weighting value for the attribute $v\in\mathcal{D}_{S}\cup\mathcal{D}_{T}$ is defined as
$\alpha(v)=\textit{Importance}(v)+\textit{Collision}(v).$ (6)
This weighting value accounts for the relative significance of each attribute in the evaluation process and is integrated into Eq. 7.
4.2 Deriving the Ideal Editing in CLIP Space
Based on the extracted source and target attributes, we describe how to derive a modification vector that satisfies i) modification into the target text and ii) preservation of the source image.
Criterion 1: Does the edited image follow the target text?
To determine if the edited image faithfully follows the given target text, it is intuitive to employ a classifier that classifies the image into either source or target. Classification into source indicates the edited image is still closer to the source image, while classification into target means that the image is sufficiently modified. The classifier is derived by a hyperplane that separates the source $\mathcal{D}_{S}$ and target $\mathcal{D}_{T}$.
Formally, the classifier function $g(x)=\mathbf{w}^{T}x+b$ assigns a CLIP embedding $x$ to the target class if $g(x)>0$, or to the source class if $g(x)<0$. Therefore, $E(I_{\text{edit}})$ should satisfy $\mathbf{w}^{T}E(I_{\text{edit}})+b>0$ if the edited image faithfully follows the target text. Here, the optimization form of the classifier function $g(x)$ is
$\min_{\mathbf{w},b}\frac{1}{2}\|\mathbf{w}\|^{2}+C\sum_{i=1}^{N}\alpha(v_{i})\cdot\max(0,1-y_{i}(\mathbf{w}\cdot v_{i}+b)),$ (7)
where $y_{i}=+1$ for $v_{i}\in\mathcal{D}_{T}$ and $y_{i}=-1$ for $v_{i}\in\mathcal{D}_{S}$. Here, the precomputed weighting values $\alpha(\cdot)$ in Eq. 6 are integrated into the optimization process.
Criterion 2: Does the edited image preserve the source image?
However, it is equally important to consider the second criterion, source image preservation. Therefore, the CLIP representation of modification should inflict minimum modification to the original embedding $E(I_{\text{src}})$, while being classified as a target class. Formally, we derive the modification $\mathbf{v}$ in CLIP space that satisfies
$\underset{\mathbf{v}}{\text{min}}\|\mathbf{v}\|\quad\text{subject to}\quad\mathbf{w}^{T}\left(E\left(I_{\text{src}}\right)+\mathbf{v}\right)+b>0.$ (8)
Derivation of $\mathbf{v}$
To minimize $\|\mathbf{v}\|$, $\mathbf{v}$ should be in the direction that most efficiently increases the classifier output $g(\cdot)$. This direction is given by the gradient of the classifier function concerning $\mathbf{x}$, which is $\mathbf{w}=\nabla g(\mathbf{x})$. Therefore, the optimal $\mathbf{v}$ is $c\mathbf{w}$ where $c>0$ is a scalar to be determined. Substituting $\mathbf{v}=c\mathbf{w}$ to Eq. 8, we derive the following equation
$c>\frac{-(\mathbf{w}^{\top}I_{\text{src}}+b)}{\|\mathbf{w}\|^{2}}.$ (9)
In order to solve for $c$ that minimizes $\|\mathbf{v}\|$, we derive the smallest $c$ satisfying the condition as $c_{\text{min}}=\frac{-(\mathbf{w}^{\top}I_{\text{src}}+b)}{\|\mathbf{w}\|^{2}}$. Then the modification vector $\mathbf{v}$ is derived as
$\mathbf{v}=c_{\text{min}}\mathbf{w}=\frac{-(\mathbf{w}^{\top}I_{\text{src}}+b)}{\|\mathbf{w}\|^{2}}\mathbf{w}.$ (10)
Finally, AugCLIP is computed by substituting $\mathbf{v}$ to Eq. 3.
Table 1: Benchmark Datasets and Editing Models. Benchmark dataset type that corresponds to different editing scenarios. For each dataset, we use different types of editing models that are tailored for each scenario.
$\dagger$ Kwon et al. (2022); Kim et al. (2022); Kim & Ye (2021); Patashnik et al. (2021)
$\ddagger$ Hertz et al. (2023); Couairon et al. (2022); Brooks et al. (2022); Hertz et al. (2022); Kawar et al. (2022)
$\ast$ Li et al. (2024a); Kumari et al. (2023); Wei et al. (2023)
CelebA EditVal TEdBench MagicBrush DreamBooth
Dataset Types Facial Attribute General Object Object Centric Local Region Editing Personalized Generation
Editing Models $\dagger$ $\ddagger$ $\ddagger$ $\ddagger$ $\ast$
Dataset Size 50 648 100 1053 50
5 Experiments
Implementation Details.
For our experiments, we employ a pre-trained CLIP-ViT 16/B model for CLIP-based metrics. Source and target attributes are generated using GPT-4V (OpenAI, 2023). Further details on prompting the source and target attributes are in the Appendix.
Baseline Metrics.
We compare AugCLIP with DINO similarity, LPIPS, L2, Segmentation Consistency (SC), Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$), CLIP-I, and CLIP-T. The CLIP variants are summarized as follows:
$\text{CLIP}_{\text{dir}}=\texttt{cs}\left(E(I_{\text{edit}})-E(I_{\text{src}}),E(T_{\text{trg}})-E(T_{\text{src}})\right)$
[…]
Note that FID measures the distance between two distributions of image features, this metric cannot be measured in a sample-wise manner. Therefore, FID cannot be tested in terms of alignment with human evaluation or preference test on ground truth samples.
Benchmark Datasets and Editing Models.
We evaluate AugCLIP and existing metrics across several text-guided image editing benchmarks, including TEdBench (Kawar et al., 2022), EditVal (Basu et al., 2023), MagicBrush (Zhang et al., 2024), DreamBooth (Ruiz et al., 2023), and CelebA Liu et al. (2015). Each benchmark dataset represents varying editing scenarios as in Tab. 1.
For each benchmark dataset, we employ multiple editing models (Kwon et al., 2022; Kim et al., 2022; Kim & Ye, 2021; Patashnik et al., 2021; Hertz et al., 2023; Couairon et al., 2022; Brooks et al., 2022; Hertz et al., 2022; Li et al., 2024a; Kumari et al., 2023; Wei et al., 2023) as specified in Tab. 1. For CelebA, which require facial attribute editing, we employ domain-specific editing models pre-trained on facial dataset. For EditVal, TEdBench, and MagicBrush, we employ generic-purpose editing models. For DreamBooth, we use presonalized generation models. More details on benchmark datasets are provided in the Appendix.
5.1 Correlation with Human Judgment
User Study
To evaluate the effectiveness of AugCLIP compared to existing metrics, we conduct a Two-Alternative Forced Choice survey on human evaluators to collect preferences over diverse edited images. The alignment score ${\bm{s}}_{\text{2AFC}}$ measures if each metric prefers the same option as human evaluators. The survey was conducted using Amazon Mechanical Turk (AMT), over three benchmark datasets, encompassing a total of 105 participants as detailed in Tab. 2. The survey questionnaire is structured as described in the Appendix, providing clear criteria to consider source image preservation and target text alignment.
AugCLIP aligns better with human judgments.
As shown in Tab. 3, AugCLIP, the context-aware metric that balances preservation and modification, is superior over all previous metrics across diverse benchmark datasets. A high level of alignment between preservation-centric (L2, LPIPS, DINO, SC, CLIP-I) and human judgment is expected because these benchmarks do not significantly deviate from the source image.
Table 2: User Study Statistics.
CelebA EditVal DreamBooth
Survey questions 39 35 37
Participants 45 30 30
Table 3: Correlation with Human Judgment. Human alignment score ${\bm{s}}_{\text{2AFC}}$ ($\uparrow$), scaled between 0 to 1.
Metrics Presv. Modif. CelebA EditVal DreamBooth
L2 0.653 0.348 0.464
LPIPS 0.465 0.360 0.286
DINO 0.574 0.348 0.286
SC 0.752 0.764 0.571
CLIP-I 0.848 0.730 0.857
CLIP-T 0.491 0.399 0.321
[…]
AugCLIP 0.883 0.831 0.857
5.2 Ground Truth Editing Selection Test
Creating Triplet Dataset
[…]
AugCLIP is superior in balancing preservation and modification.
In Tab. 4, AugCLIP is on par with or even better than baseline metrics over all datasets. This emphasizes the balanced evaluation criteria of AugCLIP, which previous metrics fail to abide by. Segment Consistency (SC) excels in MagicBrush since the dataset requires editing a small local area of the image.
Table 4: Ground Truth Selection Test. Accuracy, $\textbf{Acc}_{\text{both}}$ ($\uparrow$), of assigning higher scores to ground truth images over excessively preserved and modified images.
Metrics Presv. Modif. TEdBench MagicBrush
L2 0.310 0.002
LPIPS 0.090 0.000
DINO 0.280 0.008
SC 0.420 0.936
CLIP-I 0.281 0.810
CLIP-T 0.312 0.260
[…]
AugCLIP 0.570 0.889
6 Analysis
[…]
AugCLIP generalizes to diverse editing scenarios.
[…]
Table 5: Human Alignment Score ${\bm{s}}_{\text{2AFC}}$ on Various Text-guided Image Editing Scenarios.
Pos. Add Obj. repl. Alter Parts Background Texture Color Action Style
[…]
AugCLIP 1.0 0.75 0.838 1.0 0.742 1.0 1.0 0.647
[…]
Table 6: Effect of Augmenting Attributes.
CelebA EditVal DreamBooth TEdBench MagicBrush
[…]
$\quad+\mathcal{D}_{S}$ 0.816 0.629 0.357 0.400 0.429
$\quad+\mathcal{D}_{T}$ 0.819 0.708 0.536 0.420 0.533
$\quad+\mathcal{D}_{S}\cup\mathcal{D}_{T}$ 0.816 0.607 0.536 0.440 0.402
AugCLIP 0.883 0.831 0.857 0.570 0.889
Weighting Strategy for Hyperplane
[…]
Table 7: Effect of Weighting Strategy.
CelebA EditVal DreamBooth TEdBench MagicBrush
Unweighted 0.849 0.787 0.786 0.400 0.830
Weighted 0.883 0.831 0.857 0.570 0.889
Length and Number of Attributes
In Tab. 8, we compare short (under 5 words) and long (over 5 words), as well as cases with a fixed number (10, 20, or 30) of attributes. We observe that using short attributes tends to outperform using long ones, likely because they focus on a single visual aspect, avoiding the potential entanglement of multiple aspects in a single attribute. Moreover, the number of attributes impacts the performance depending on the scene complexity of the benchmark dataset, as complex scenarios require a higher number of attributes to fully capture the scene’s details. Therefore, allowing flexibility in the number of attributes, rather than imposing a strict limit, yields the best overall performance across all configurations.
Table 8: Effect of Length and Number of Attributes.
Length Number CelebA EditVal DreamBooth TEdBench MagicBrush
short 10 0.870 0.719 0.857 0.540 0.889
short 20 0.829 0.809 0.821 0.540 0.868
short 30 0.829 0.764 0.714 0.570 0.863
long 30 0.843 0.697 0.750 0.530 0.862
short - 0.883 0.831 0.857 0.570 0.889
Description Randomness
Using GPT-4V for description extraction introduces slight randomness across different seeds. Tests with five seeds show low variance ($0.0197$) in human alignment score (${\bm{s}}_{\text{2AFC}}$), indicating that the description generation process does not affect the scoring. We also demonstrate in the Appendix that extracted attributes with different seeds do not vary.
We provide analysis on the modification vector $\mathbf{v}$ and choice of classifier function, discussion on limitations, description prompting, and extensive qualitative results in the Appendix.
Limitations and Discussions
Our method, AugCLIP, adds additional computation time compared to $\mathrm{CLIP}_{\text{dir}}$. It requires extracting attributes via MLLM, and then fitting a hyperplane to distinguish source and target attributes. This process takes around 12.3 seconds for attribute generation and 0.15 seconds for score computation. We will provide extracted attributes for all benchmark datasets to reduce this overhead, facilitating the evaluation process for future usage.
Conclusion
We present AugCLIP, a novel evaluation metric for text-guided image editing that balances source image preservation and target text modification. By leveraging a multi-modal language model to extract visual attributes and finding a separating hyperplane, AugCLIP estimates a representation of an ideal edited image that closely matches human preferences. Experiments on five benchmark datasets show that AugCLIP outperforms existing metrics, especially in challenging tasks, offering a more reliable tool for evaluating edits while preserving core attributes.
Appendix A Discussions
A.1 Intuitive Difference Between AugCLIP and Directional CLIP Similarity
[…]
In real-world evaluation cases, this is an important property since some tasks might require a small amount of editing that transforms only part of the image whereas some tasks focus on editing the whole image with large modifications. Evaluation metrics should be flexibly applicable to all such cases, deciding the evaluation standard to focus on preservation or modification on a case-by-case basis. However, regardless of the editing context requiring different modification or preservation levels, existing metrics blindly apply the same standard that overly focuses on preservation or modification.
Figure 5: Difference between $\mathrm{CLIP}_{\text{dir}}$ and AugCLIP. The red line indicates the evaluation standard and the black line indicates the change in image from source to target. Both $\mathrm{CLIP}_{\text{dir}}$ and AugCLIP measure the quality of the edited image according to the corresponding red lines. In (b), the green and yellow circles indicate the distribution of target and source attributes, respectively, and dotted black lines indicate the evaluation standard of $\mathrm{CLIP}_{\text{dir}}$.
A.2 Randomness of Descriptions
As discussed in Sec. 6, the randomness of extracted descriptions does not impact the alignment with human judgment (merely $0.0197$ in variance). In addition to this observation, we provide description samples for the target text, varying the seed over five configurations in Fig. 6, 7. The examples show that generated descriptions have overlapping semantics across different seeds. For example, Fig. 7(a) across five seeds describe black and white stripes, with small textual differences but identical in terms of semantics. This is observed across all four examples of the target text, proving that randomness in the description extraction process does not create semantically distinct samples, thus the evaluation results of AugCLIP are robust. the difference across random seeds is almost negligible since source descriptions are directly extracted from the source image caption.
(a) Target text: Change a standing dog into a sitting dog.
(b) Target text: Change a closed box into an open box.
Figure 6: Target Descriptions Generated with Five Random Seeds.
(a) Target text: Change a horse into a zebra.
(b) Target text: Make a cat yawn.
Figure 7: Target Descriptions Generated with Five Random Seeds.
A.3 Optimization Objective
In Tab. 9, we test three variants of optimization objective for deriving the classifier function $g(x)=\mathbf{w}^{T}x+b$. Since source and target descriptions encoded into CLIP are separable by a simple linear function, we set the hyperplane as $\mathbf{w}^{T}x+b=0$. Moreover, in optimizing the parameters $\mathbf{w}$ and $b$, we select the hinge loss objective with L2 regularization, namely SVM objective.
Reason for Choosing SVM Objective.
We have emonstrate in Tab. 9, that this SVM objective is the best option. Moreover, the reasons for choosing this objective are threefold. First, the SVM objective is compatible with various source and target cases. Since SVM allows for misclassification to some extent with the slack variable, it extends to editing cases where the source image and target text do not show distinct discrepancies. Second, support vectors maximize the margin, which is defined as the distance between the hyperplane and the closest support vectors on either source and target descriptions. Third, the influence of data points far from the margin is reduced, since only the closest points (support vectors) determine the decision boundary. This helps avoid bias caused by outlying data points, ensuring the hyperplane is not skewed toward either the preservation or modification side.
Comparing Optimization Objectives.
In order to test if choosing other hyperplane optimization objective impacts the level of alignment with human judgment, we compare latent discriminant analysis (LDA), logistic regression, and linear SVM objective in finding a separating hyperplane. More specifically, the objective functions are as follows,
LDA $:\frac{\text{det}(N_{S}(\mu_{S}-\mu)(\mu_{S}-\mu)^{T}+N_{T}(\mu_{T}-\mu)(\mu_{T}-\mu)^{T})}{\text{det}(\sum_{{\bm{s}}_{i}\in\mathcal{D}_{S}}({\bm{s}}_{i}-\mu_{S})({\bm{s}}_{i}-\mu_{S})^{T}+\sum_{{\bm{t}}_{j}\in\mathcal{D}_{T}}({\bm{t}}_{j}-\mu_{T})({\bm{t}}_{j}-\mu_{T})^{T})}$
Logistic $:-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\sigma(g(x_{i})))+(1-y_{i})\log(1-\sigma(g(x_{i})))\right]$
SVM $:\frac{1}{2}\|\mathbf{w}\|^{2}+C\sum_{i=1}^{N}\max(0,1-y_{i}\cdot g(x_{i}))$
where the optimization targets to find $\mathbf{w}$ and $b$ that minimizes the given objective functions. A total of $N$ pairs of data points $(x,y)$ is employed in the optimization process, where $x$ in these objective functions signifies the CLIP-encoded source or target attributes with the corresponding label $y\in\{-1,1\}$.
[…]
Table 9: Comparison on difference optimization function for the classifier.
CelebA EditVal DreamBooth TEdBench MagicBrush Average Misc.
LDA 0.884 0.827 0.821 0.545 0.863 3.37%
Logistic 0.849 0.830 0.821 0.550 0.866 1.38%
Linear SVM 0.883 0.831 0.857 0.570 0.889 1.35%
Misclassification Rate.
Our metric AugCLIP first encodes the source and target attributes into CLIP space, denoted as $\mathcal{D}_{S}$ and $\mathcal{D}_{T}$ respectively. We then employ a linear function $g(x)=\mathbf{W}^{T}x+b$ to estimate a decision boundary that separates the source and target distribution. The average percentage of source and target attributes that are wrongly classified by the hyperplane across five benchmark datasets is reported in Tab. 9, denoted as ‘Average Misc.’. We observe that simple linear decision function $g(x)$ shows a small misclassification rate, $1.35\%$, which signifies its ability to separate the source and target distributions. Specifically, linear SVM achieves the lowest misclassification rate, successfully distinguishing between source and target attributes. Given that the source image and target text may share visual similarities, the extracted source and target attributes cannot always be perfectly separable by a linear hyperplane (e.g., when editing an orange to a tangerine, both the source and edited images share a round shape.). In such cases, these attributes are closely positioned in the embedding space and do not require complete separation. SVM’s ability to manage overlapping factors more flexibly allows it to find a more accurate hyperplane, leading to superior performance.
UMAP Visualization.
Additionally, we visualize that CLIP features of source and target attributes can be separated by a linear hyperplane in 2D projected space using UMAP (McInnes et al., 2018) in Fig. 8, in which randomly chosen subset of TEdBench samples are plotted. ‘S’ represents source attributes encoded into CLIP, while ‘T’ represents the target attributes. The line signifies the linear hyperplane $g(x)=\mathbf{W}^{T}x+b=0$ that separates the two classes, which are source and target. In AugCLIP, the linear hyperplane $g(x)=0$ is $d$-dimensional following the original dimension of CLIP, but to visualize in Fig. 8, the dimension shrinks into $d=2$ by UMAP fitting.
Figure 8: Visualization of Source and Target Attributes in 2D Space. $S$ indicates the source attributes encoded into CLIP space and $T$ indicates the target attributes. The separating hyperplane is a linear function $\mathbf{W}^{T}x+b=0$.
A.4 Analyzing the Modification Vector
The goal of text-guided image editing is to apply a necessary transformation to the source image, preserving the original property as much as possible. Our evaluation process follows this protocol by estimating the modification vector $\mathbf{v}$ as a minimum modification that makes the source image look like a target text, ensuring that essential source attributes remain unchanged but the resulting edited image resembles the target text. In this section, we qualitatively analyze the effect of $\mathbf{v}$ on source and target attributes.
Figure 9: Effect of the Modification Vector $\mathbf{v}$ on Source Attributes. The source attributes are ranked based on the magnitude of change induced by the modification vector $\mathbf{v}$. Attributes at the top of the list exhibit the most significant adjustments toward the target text, indicating that these characteristics are considered important source attributes that should be modified in the given editing context. Conversely, attributes lower down the list are determined to remain intact in an ideal editing.
Effect of the Modification Vector $\mathbf{v}$ on Source Attributes
Source attributes that need to be preserved should not be affected by the modification vector $\mathbf{v}$, while those requiring adjustment according to the target text should be altered. To demonstrate that $\mathbf{v}$ drives significant changes in the attributes requiring modification while minimally impacting those that should remain intact, we analyze several cases using the TEdBench and EditVal datasets, as shown in Fig. 9. We measure the difference in cosine similarity between each source attribute ${\bm{s}}_{i}$ and both the source image and the ideally edited image. Specifically, we calculate the increase in similarity between the ideal edited image $E(I_{\text{src}})+\mathbf{v}$, and the source image $E(I_{\text{src}})$ as
$\Delta{\bm{s}}_{i}=\texttt{cs}\Big{(}E(I_{\text{src}})+\mathbf{v},{\bm{s}}_{i}\Big{)}-\texttt{cs}\Big{(}E(I_{\text{src}}),{\bm{s}}_{i}\Big{)},$ (11)
where ${\bm{s}}_{i}$ indicate the CLIP-encoded source attribute in $\mathcal{D}_{S}$. As demonstrated in Fig. 9, source attributes that need to be preserved exhibit only small changes in similarity (small $\Delta\mathbf{s}_{i}$), while attributes that need modification (large $\Delta\mathbf{s}_{i}$) show a significant change in similarity.
Figure 10: Effect of the Modification Vector $\mathbf{v}$ on Target Attributes. The target attributes are ranked based on the magnitude of change induced by the modification vector $\mathbf{v}$. Attributes at the top of the list are determined by AugCLIP standard to be key to modification aspects. Attributes lower on the list exhibit smaller adjustments, which are deemed less important by AugCLIP standard.
Effect of the Modification Vector $\mathbf{v}$ on Target Attributes
We demonstrate how target attributes are affected by the modification vector $\mathbf{v}$. To demonstrate that $\mathbf{v}$ causes significant changes in the attributes that are central to make the image resemble the target text while having minimal impact on the rather peripheral attributes, we analyze several cases using the EditVal dataset, as shown in Fig. 10. The difference in cosine similarity between the source image and the ideally edited image is measured for each target attribute ${\bm{t}}$. Similar to the previous paragraph, we compare the increase of target attribute ${\bm{t}}_{i}$ in the ideal edited image, $E(I_{\text{src}})+\mathbf{v}$, compared to the source image $E(I_{\text{src}})$ as
$\Delta{\bm{t}}_{i}=\texttt{cs}\Big{(}E(I_{\text{src}})+\mathbf{v},{\bm{t}}_{i}\Big{)}-\texttt{cs}\Big{(}E(I_{\text{src}}),{\bm{t}}_{i}\Big{)},$ (12)
where ${\bm{t}}_{i}$ indicate the CLIP-encoded target attribute in $\mathcal{D}_{T}$. As illustrated in Fig. 10, attributes essential for matching the target text display a significant increase in similarity, while secondary attributes experience only minor changes.
Appendix B Algorithm
We provide the algorithm of AugCLIP in Python code style in the following block.
1# Step 0: Get CLIP features
2src_img_feat, tgt_img_feat = CLIP(src_img), CLIP(tgt_img)
3src_text_feat, tgt_text_feat = CLIP(src_text), CLIP(tgt_text)
4src_desc_feat, tgt_desc_feat = CLIP(src_desc), CLIP(tgt_desc)
5
6# Step 1: Compute importance weighting for each desc
7src_dist = [src_img_feat, src_desc_feat]
8tgt_dist = [tgt_img_feat, tgt_desc_feat]
9
10src_weight, tgt_weight = compute_weight(src_dist, tgt_dist)
11weight = [src_weight, tgt_weight]
12
13# Step 2: Fit the classifier model
14X = [src_dist, tgt_dist]
15y = [-1] * src_dist.shape[0] + [1] * tgt_dist.shape[0]
16svc_classifier.fit(X, y, sample_weight=weight)
17
18# Step 3: Retrieve the hyperplane parameters
19w = svc_classifier.coef_ # Hyperplane coefficients
20b = svc_classifier.intercept_ # Hyperplane intercept
21
22# Step 4: Compute the modification vector v by calculating the projection of src_img_feat onto the hyperplane
23numerator = -(np.dot(w.T, src_img_feat) + b)
24denominator = np.linalg.norm(w)**2
25v = (numerator / denominator) * w
26
27# Step 5: Calculate the alignment score
28score = cosine_similarity(src_img_feat + v, tgt_img_feat)
Appendix C Evaluation Details
C.1 Assets
Table 10: Assets Employed in Our Experiments. List of pre-trained models, benchmark datasets, and metrics employed in this paper.
Category Asset URL
Benchmarks CelebA (Liu et al., 2015) https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
TedBench (Kawar et al., 2022) https://github.com/imagic-editing/imagic-editing.github.io/tree/main/tedbench
EditVal (Basu et al., 2023) https://github.com/deep-ml-research/editval_code
DreamBooth (Ruiz et al., 2023) https://github.com/google/dreambooth
MagicBrush (Zhang et al., 2024) https://github.com/OSU-NLP-Group/MagicBrush
Editing Models TEdBench EditVal MagicBrush InstructPix2Pix (Brooks et al., 2022) https://github.com/timothybrooks/instruct-pix2pix
DiffEdit (Couairon et al., 2022) https://github.com/Xiang-cd/DiffEdit-stable-diffusion.git
Prompt-to-Prompt (Hertz et al., 2022) https://github.com/google/prompt-to-prompt.git
DDS (Hertz et al., 2023) https://github.com/google/prompt-to-prompt/blob/main/DDS_zeroshot.ipynb
Imagic (Kawar et al., 2022) https://github.com/huggingface/diffusers/tree/main/examples/community#imagic-stable-diffusion
CelebA DiffusionCLIP (Kim & Ye, 2021) https://github.com/gwang-kim/DiffusionCLIP
Multi2One (Kim et al., 2022) https://github.com/akatigre/multi2one
Asyrp (Kwon et al., 2022) https://github.com/kwonminki/Asyrp_official
StyleCLIP (Patashnik et al., 2021) https://github.com/orpatashnik/StyleCLIP
DreamBooth Custom Diffusion  (Kumari et al., 2023) https://github.com/adobe-research/custom-diffusion
BLIP-Diffusion (Li et al., 2024a) https://github.com/salesforce/LAVIS/tree/main/projects/blip-diffusion
ELITE (Wei et al., 2023) https://github.com/csyxwei/ELITE
Metrics CLIP (Radford et al., 2021) https://github.com/openai/CLIP
LPIPS (Zhang et al., 2018) https://pypi.org/project/lpips/
SC (Kim & Ye, 2021) Implemented by the Authors
C.2 Source Description Generation
As described in Sec. 4.1, we prompt GPT-4V with the source image to first generate a detailed caption. Specifically, the prompt orders to focus on the visual and objective aspects of the image, disregarding any opinions or non-visual descriptions. With the generated caption, we provide an example of another source image caption and its parsed list of visual characteristics. With these examples, GPT-4V parses the source image caption into a Python list of visual characteristics. The example prompting result is shown as follows.
User [Uncaptioned image] Describe the image in detail. Do not include opinions or non-visual aspects. Make the descriptions easy and intuitive.
GPT-4V The image shows a person sitting on a low ledge. He is wearing an orange T-shirt and tan shorts. His legs are spread apart, and he appears to be focused on his phone, which he is holding with both hands. He has a gray cap on his head, and his athletic shoes are orange with gray accents. His posture is slightly hunched over as he engages with his device.
User You are a helpful text-generation assistant. Given detailed textual descriptions about an image, your goal is to parse into specific visual attributes. EXAMPLE INPUT: “The image features a large, dark-colored dog standing in a grassy field. The dog appears to be alert and attentive, possibly observing its surroundings. The grass is lush and green, providing a natural backdrop for the dog.” EXAMPLE OUTPUT: [“A dog is large”, “A dog is dark-colored”, “A dog is standing”, “A dog is standing on a grassy field”, “A dog is alert”, “A dog is attentive”, “A dog is observing its surroundings”, “A grass is lush”, “A grass is green”, “A grass is providing a natural backdrop”]   Give me a Python list of visual characteristics included in the following sentences:
INPUT: “The image shows a person sitting on a low ledge. He is wearing an orange T-shirt and tan shorts. His legs are spread apart, and he appears to be focused on his phone, which he is holding with both hands. He has a gray cap on his head, and his athletic shoes are orange with gray accents. His posture is slightly hunched over as he engages with his device.”
GPT-4V [“He is sitting on a low ledge”, “He is wearing an orange t-shirt”, “He is wearing tan shorts”, “He has his legs spread apart”, “He is focused on his phone”, “He is holding the phone with both hands”, “He has a gray cap on his head”, “He has athletic shoes”, “He has orange shoes with gray accents”, “He is slightly hunched over”]
C.3 Target Description Generation
The descriptions for the target text are generated by providing the target text and the source image caption. The prompt instructs the model to avoid repetition of similar visual attributes. The example depicts a case of changing a sitting person into a standing posture. The output is structured into a Python list of descriptions, starting with “Standing person”. From the target text “Change a man into Ironman,” the generated list of attributes includes descriptions of the suit, weapons, and armor. These also depict that Ironman takes heroic poses and appears in intense battles. The example prompting result is shown as follows.
User You are a helpful text-generation assistant. Given a detailed textual description of an image, your goal is to parse it into specific visual attributes. If the visual attributes with similar meanings are repeated, only use one. For example, “A dog is large” and “A dog is big” are similar, so only use one. EXAMPLE INPUT: In order to make a person into a standing posture, what visual attributes of the image should be changed? Answer in the format of “Standing person is ATTRIBUTE”. EXAMPLE OUTPUT: [“Standing person has straight legs”, “Standing person is upright”, “Standing person is on their feet”, “Standing person is tall”]   INPUT: In order to make a person into Ironman, what visual attributes of the image should be changed? Answer in the format of “Ironman is ATTRIBUTE”.
GPT-4V [“Ironman has powered exoskeleton suit”, “Ironman has a red and gold color scheme”, “Ironman has Arc reactor on chest”, “Ironman wears metallic armor”, “Ironman wears helmet with facial display”, “Ironman has repulsor technology on hands”, “Ironman has jet boots for flight”, “Ironman has advanced weapon systems”, “Ironman has a strong and muscular build”, “Ironman has heroic poses”, “Ironman appears in intense battles” ]
C.4 User Study Details
Figure 11: User Study Details. The figure shows the instructions and example questions of our user study.
As existing text-guided image editing models do not guarantee adequate visual quality, we manually select the images that show sufficient change in the image to conduct a user study. Each participant is provided with a source image, its corresponding target text, and two variants of edited images. Then, human evaluators are instructed to choose the image with better editing quality. As shown in Fig. 11, clear guidelines are provided to instruct the participants to evaluate the images based on both the preservation of the source image and the modifications toward the target text.
C.5 Benchmark Datasets
TEdBench comprises $100$ pairs of source image and target text. It focuses on specific settings where the source image has a single object at the center, and the corresponding target text only modifies some attributes of that object.
EditVal contains $648$ image-text pairs that cover $13$ different types of edits, including object addition, object replacement, and size modification. Since it has such complicated editing scenarios, the models that we use for editing could not properly edit the majority of the cases, leaving almost no samples with enough quality for user study. Therefore, we use the subset of EditVal, which encompasses eight editing types that show adequate modification for proper evaluation.
MagicBrush is a benchmark specifically designed to evaluate sequential editing tasks, where iterative modifications are made to different parts of the source image.
Dreambooth enables the modification of specific instances within the source image by providing corresponding masks along with image-text pairs; however, since typical editing models do not utilize masks as input, we only consider the image-text pairs in our evaluation.
CelebA dataset consists of $50$ image-text pairs that guide changes specific to facial attributes. We create target texts by swapping attributes of human faces.
Appendix D Additional Results
D.1 Combination of Preservation and Modification Centric Metrics
As discussed in Sec. 3.1, combining modification-centric metric (CLIP-T) with existing preservation-centric metrics (DINO similarity, Segment Consistency, CLIP-I) shows negligible improvement or rather deteriorates in terms of alignment with human judgment and ground truth selection test. Due to the spatial constraint, we have shown two of the datasets, EditVal and CelebA, in the main paper. In Fig. 12, we demonstrate the results of the other three datasets. Notably, the combination of CLIP-I and CLIP-T shows improvement in the TEdBench dataset. Since TEdBench is the dataset of the simplest setting, where source images are highly object-centric and the target text instructs relatively simple modification, a simple strategy of combining these two metrics could be a viable option for evaluation. However, as shown in most editing cases, such a simplistic combination approach fails to show large improvement, underscoring the need for a metric tailored for text-guided image editing.
Figure 12: Interpolation of Preservation- and Modification-Centric Metrics.
D.2 Additional Result on Augmenting Directional CLIP Similarity
[…]
Weighting CelebA EditVal DreamBooth TEdBench MagicBrush
$\mathrm{CLIP}_{\text{dir}}$ 0.673 0.697 0.357 0.350 0.601
+ src attr. 0.816 0.629 0.357 0.400 0.429
0.796 0.725 0.464 0.440 0.523
+ trg attr. 0.819 0.708 0.536 0.420 0.533
0.734 0.513 0.464 0.450 0.443
+ src & trg attr. 0.816 0.607 0.536 0.440 0.402
0.636 0.600 0.536 0.500 0.407
AugCLIP 0.883 0.831 0.857 0.570 0.889
In Tab. 6, we demonstrate that the simple strategy of directly augmenting $\mathrm{CLIP}_{\text{dir}}$ with the source and target attributes fails to outperform AugCLIP. Additionally, we show the effect of applying the weighting strategy () of Eq. 6 when aggregating attributes into $\mathrm{CLIP}_{\text{dir}}$ in Tab. 11. Note that augmentation of $\mathrm{CLIP}_{\text{dir}}$ without weighting () is already reported in Tab. 6.
Formally, for augmenting the source text $T_{\text{src}}$ with source attributes in $\mathcal{D}_{S}$, directional CLIP similarity is redefined as
$\texttt{cs}\Big{(}E(I_{\text{edit}})-E(I_{\text{src}}),E(T_{\text{trg}})-\mathbb{E}_{\mathbf{s}_{i}\in\mathcal{D}_{S}}(s_{i})\Big{)},$ (13)
where $\mathbb{E}$ means expectation. Using the weighting strategy with $\alpha$ defined in Eq. 6, $\mathrm{CLIP}_{\text{dir}}$ is reformulated as
$\texttt{cs}\left(E(I_{\text{edit}})-E(I_{\text{src}}),E(T_{\text{trg}})-\mathbb{E}_{\mathbf{s}_{i}\in\mathcal{D}_{S}}\left(\alpha(\mathbf{s}_{i})\cdot\mathbf{s}_{i}\right)\right).$ (14)
The same formulation applies for the target text $T_{\text{trg}}$ as well.
Across all configurations, with and without weighting, AugCLIP outperforms $\mathrm{CLIP}_{\text{dir}}$ in terms of alignment with human judgment and ground truth selection test accuracy. This emphasizes that our metric, AugCLIP, notably well-performs compared to $\mathrm{CLIP}_{\text{dir}}$.
D.3 Comparison with GPT-4V
[…]
CelebA EditVal DreamBooth TEdBench MagicBrush
GPT-4V 0.876 0.933 0.821 0.620 0.703
AugCLIP 0.883 0.831 0.857 0.570 0.889
Recently, GPT-4V (OpenAI, 2023) has been employed in evaluating various tasks, including text-guided image editing, text-to-image generation, and image quality assessment. Since GPT-4V is one of the best-performing multi-modal large language models, we test the ability of GPT-4V’s effectiveness in evaluating the quality of text-guided edited images. For evaluation, we use the following prompt: “Given a source image and two edited images, you should choose a better edited one based on the source and target text. Source text describes the source image, and target text describes the editing. A well-edited image should preserve the essence of the source image while following the target text.”
As shown in Tab. 12, GPT-4V outperforms AugCLIP in tasks such as EditVal and TEdBench, which involve simple edits like modifying a single object’s attribute. In contrast, our proposed metric, AugCLIP, effectively captures minor differences by augmenting attributes of the source image and target text and shows better performance in other benchmarks with complex scenarios. This finding is consistent with prior research (Zhang et al., 2023), which suggests that GPT-4V struggles to differentiate between images with subtle differences.
D.4 Additional Examples on Problem 1 of Directional CLIP Similarity
[…]
D.5 Additional Examples on Problem 2 of Directional CLIP Similarity
[…]
To assess directional CLIP similarity’s capability to focus on the image regions modified following the target text rather than peripheral or unchanged regions, we use the relevancy map (Chefer et al., 2021), ${\bm{R}}$. The relevancy map visualizes the transformer’s attention on an image corresponding to a given text depending on their cosine similarity. Specifically, for an image $I\in\mathbb{R}^{h\times w}$ and text $T$, the relevancy map is computed as
${\bm{R}}(I;T)=\nabla_{\bm{A}}\texttt{cs}(E(I),E(T);{\bm{A}})\odot{\bm{A}}\in\mathbb{R}^{h\times w},$
where ${\bm{A}}$ represents the attention scores of the CLIP visual encoder and $\odot$ denotes the Hadamard product. To visualize the relevancy map of $\mathrm{CLIP}_{\text{dir}}$, which is a cosine similarity between $\Delta I$ and $\Delta T$, we subtract the two relevancy maps as
${\bm{R}}(\Delta I;\Delta T)={\bm{R}}(I_{\text{edit}};\Delta T)-{\bm{R}}(I_{\text{src}};\Delta T).$
Fig. 3(b) and Fig. 14 illustrate the relevancy maps of $\mathrm{CLIP}_{\text{dir}}$ across multiple cases and their improvement achieved by incorporating manually annotated visual descriptions. Unlike $\mathrm{CLIP}_{\text{dir}}$, AugCLIP measures the cosine similarity between the estimated well-edited image and the edited image, rather than between an image and text. As a result, the relevancy map, which requires a direct comparison of the image and text, cannot be applied to AugCLIP.
Appendix E Qualitative Results
We present qualitative samples of the 2AFC Test, as reported in Tab. 3, using the CelebA, EditVal, and DreamBooth datasets. For each dataset, we randomly select triplets consisting of a source image, target text, and edited images to demonstrate how AugCLIP consistently assigns higher scores to the edited image preferred by human evaluators. The preferred image, highlighted with a red box, appears in the middle. Each case represents a two-alternative forced choice (2AFC) survey, where the source image on the far left is altered into the middle and rightmost images. We observe that directional CLIP similarity often favors excessively modified images. For instance, in the second row of Fig. 15, where the target text is “high arch of the eyebrows,” directional CLIP similarity prefers an edited image that changes the gender of the source image into a man. Similarly, when the target text is “wrinkle-free skin,” directional CLIP similarity assigns a higher score to an image where the hair bangs are missing.
E.1 CelebA
Figure 15: Qualitative Results on CelebA (2AFC Test).
E.2 EditVal
Figure 16: Qualitative Results on EditVal (2AFC Test).
E.3 DreamBooth
Figure 17: Qualitative Results on DreamBooth dataset (2AFC test).
Additionally, we provide qualitative samples from the Ground Truth Selection Test, reported in Tab. 4, using the TEdBench and MagicBrush datasets (Fig. 18, 19 and Fig. 20, 21). In these cases, the ground truth image is located in the second column, the excessively preserved image in the third column, and the excessively modified image in the fourth column. Once again, we observe that directional CLIP similarity tends to prefer excessive modifications.
E.4 TEdBench
Figure 18: Qualitative Results on TEdBench (Ground Truth Selection Test).
Figure 19: Qualitative Results on TEdBench (Ground Truth Selection Test).
E.5 MagicBrush
Figure 20: Qualitative Results on MagicBrush (Ground Truth Selection Test).
Figure 21: Qualitative Results on MagicBrush (Ground Truth Selection Test).
E.6 Failure Cases of AugCLIP
Compared to directional CLIP similarity, AugCLIP shows superior alignment with human evaluation and a stronger ability to classify ground truth images. However, there are several cases where directional CLIP similarity aligns closely with human preferences. Fig. 22 illustrates examples where AugCLIP diverges from human judgment.
For instance, in the first-row example, both edited images are adequately modified from the source to resemble the target text “dog.” However, the middle image emphasizes dog-like features more prominently while the right image exhibits subtler changes. Human evaluators tend to favor the more prominently modified one. In the example of adding fruit toppings to donuts, both edited images accurately depict fruit toppings while preserving the original content. Yet, human evaluators prefer the middle image, which better retains the original donut’s color and texture. Here, preference is skewed toward better preservation.
Although the edits in these examples are well-executed in terms of balancing preservation and modification, human preferences remain inherently subjective and vary significantly from case to case. This highlights the limitation of evaluation metrics in fully capturing the nuances of human judgment.
Figure 22: Failure Cases of AugCLIP. Failure cases where directional CLIP similarity correctly assigns higher evaluation scores to images that human evaluators prefer, while AugCLIP fails to.

### v3

@@
 Extensive experiments on five benchmark datasets, encompassing a diverse range of editing scenarios, show that AugCLIP aligns remarkably well with human evaluation standards, outperforming existing metrics.
-The code will be open-sourced for community use.
-\$*\$\$*\$footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
+The code is available at https://github.com/augclip/augclip_eval.
 1 Introduction
-Building on advancements in vision-language models (Radford et al., 2021; Li et al., 2022; Geng et al., 2023), recent generative models (Kawar et al., 2022; Brooks et al., 2022; Hertz et al., 2022) have been widely utilized as creative tools for image editing via text instructions.
+Building on advancements in vision-language models [24, 19, 7], recent generative models [12, 2, 8] have been widely utilized as creative tools for image editing via text instructions.
 Text-guided image editing models enable the modification of images in response to textual guidance, ensuring that changes are aligned with the provided instructions.
@@
 Despite the remarkable advancements in editing models, there has been a lack of rigorous evaluation methods, tailored specifically for text-guided image editing.
-Consequently, most studies (Hertz et al., 2023; Basu et al., 2023; Gal et al., 2022; Kim & Ye, 2021; Brooks et al., 2022; Ruiz et al., 2023; Kocasari et al., 2022) have heavily relied on human evaluation, which provides balanced consideration of preservation and modification aspects.
-However, as it is costly and impractical for real-world applications, researchers have adapted automatic evaluation metrics (Heusel et al., 2017; Zhang et al., 2018; Caron et al., 2021; Hessel et al., 2021) originally designed for other vision tasks, such as image generation or captioning.
+Consequently, most studies [9, 1, 6, 13, 2, 26, 15] have heavily relied on human evaluation, which provides balanced consideration of preservation and modification aspects.
+However, as it is costly and impractical for real-world applications, researchers have adapted automatic evaluation metrics [11, 29, 3, 10] originally designed for other vision tasks, such as image generation or captioning.
 In text-guided editing scenarios, it is essential to assess which elements in the source to preserve and which in the target to modify into, based on the specific context of the given source image and target text (see Fig.
@@
 Even combining them fails to overcome the context blindness problem.
-Among existing metrics, directional CLIP similarity (Gal et al., 2022) is unique because it provides a method to consider both the source image and target text in the evaluation.
+Among existing metrics, directional CLIP similarity [6] is unique because it provides a method to consider both the source image and target text in the evaluation.
 Even this metric suffers from a context-blindness problem, using a fixed standard without considering the specific relationship between the source image and target text.
@@
 Preservation-Centric Metrics
-FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and L2 distance measure feature-level distance between images.
+FID [11], LPIPS [29], and L2 distance measure feature-level distance between images.
 These metrics are primarily used in image generation tasks to compare the distributions of real and generated images.
 When applied to editing scenarios, they assess the distance between the source and edited images, completely ignoring the target text guiding the editing process.
-Segmentation consistency (SC) (Kim & Ye, 2021) measures the structural change.
-Additionally, DINO similarity (DINO) and CLIP-I evaluate the semantic similarity between the source and edited images within the DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) embedding spaces, also focusing solely on preservation.
+Segmentation consistency (SC) [13] measures the structural change.
+Additionally, DINO similarity (DINO) and CLIP-I evaluate the semantic similarity between the source and edited images within the DINO [3] and CLIP [24] embedding spaces, also focusing solely on preservation.
 Modification-Centric Metrics
-To assess how well an edited image aligns with the target text, the primary approach involves using a pre-trained multi-modal model (Radford et al., 2021).
-CLIPScore (CLIP-T) (Hessel et al., 2021) is the most widely adopted metric in this category, measuring the similarity between the edited image and the target text.
+To assess how well an edited image aligns with the target text, the primary approach involves using a pre-trained multi-modal model [24].
+CLIPScore (CLIP-T) [10] is the most widely adopted metric in this category, measuring the similarity between the edited image and the target text.
 This metric completely ignores the source image that should be considered for the preservation aspect.
 Holistic Evaluation of Preservation and Modification
-Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$) (Gal et al., 2022) is unique among existing metrics as it assesses both preservation and modification aspects by evaluating directional alignment between the original and edited images concerning the source and target text.
+Directional CLIP similarity ($\text{CLIP}_{\text{dir}}$) [6] is unique among existing metrics as it assesses both preservation and modification aspects by evaluating directional alignment between the original and edited images concerning the source and target text.
 However, these metrics employ a fixed evaluation standard, disregarding the context of each editing scenario, which leads to context blindness.
@@
 3, combining preservation- and modification-centric metrics or using directional CLIP similarity fails to balance preservation and modification adequately, resulting in misalignment with human judgment.
-Metrics Used in Text-to-Image Generation
-While text-to-image generation tasks benefit from a variety of evaluation metrics utilizing pre-trained multi-modal large language models (Hu et al., 2023; Li et al., 2024b; Lu et al., 2024; Huang et al., 2023; Cho et al., 2024) or VQA models (Hu et al., 2023; Ghosh et al., 2024), these metrics focus solely on prompt faithfulness.
-In contrast, image editing scenarios require a balance of image preservation and prompt adherence, making these metrics irrelevant to the evaluation of editing models.
 3 Problems of Existing Metrics in Balancing Preservation and Modification Aspects
@@
 The results show that combining rather degrades the alignment with human judgment.
+[…]
 Existing evaluation metrics, that focus on either one of the preservation or modification aspects, face a challenge in assessing both aspects.
@@
 3.1 Combination of Preservation- and Modification-Centric Metrics
-We explore whether combining preservation- and modification-centric metrics can improve alignment with human judgments across two benchmark datasets: EditVal (Basu et al., 2023) and CelebA (Liu et al., 2015) (Results on the other three benchmark datasets are deferred to the Appendix).
+We explore whether combining preservation- and modification-centric metrics can improve alignment with human judgments across two benchmark datasets: EditVal [1] and CelebA [20] (Results on the other three benchmark datasets are deferred to the Appendix).
 To do so, we combine CLIP-T, a modification-centric metric, and preservation-centric metrics (LPIPS, DINO similarity, CLIP-I) with varying interpolation values $\gamma\in[0,1]$.
@@
 This highlights the need for a tailored metric that is contextually aware of both preservation and modification, ensuring a reliable alignment with human evaluations across diverse datasets and tasks.
-[…]
 3.2 Directional CLIP Similarity
@@
 3(a).
-Specifically, $\mathrm{CLIP}_{\text{dir}}$ metric fails on $63\%$ of the 100 editing cases in TEdBench (Kawar et al., 2022) and $35.1\%$ of the 1053 cases in MagicBrush (Zhang et al., 2024) (Details in Sec.
+Specifically, $\mathrm{CLIP}_{\text{dir}}$ metric fails on $63\%$ of the 100 editing cases in TEdBench [12] and $35.1\%$ of the 1053 cases in MagicBrush [28] (Details in Sec.
 5).
@@
 $\mathrm{CLIP}_{\text{dir}}$ often focuses on peripheral or unmodified local areas of the edited image, overlooking the specific areas that are edited by the text.
-Using the relevancy map (Chefer et al., 2021), ${\bm{R}}(\Delta I;\Delta T)$ [^1] The original relevancy map ${\bm{R}}$ utilizes an image-text pair.
+Using the relevancy map [4], ${\bm{R}}(\Delta I;\Delta T)$ [^1] The original relevancy map ${\bm{R}}$ utilizes an image-text pair.
 To adapt it to $\mathrm{CLIP}_{\text{dir}}$, we subtract two relevancy maps.
@@
 As described in Fig.
-4(a), we extract visual descriptions about attributes of the source image and target text using a state-of-the-art multi-modal large language model (MLLM), GPT-4V (OpenAI, 2023).
+4(a), we extract visual descriptions about attributes of the source image and target text using a state-of-the-art multi-modal large language model (MLLM), GPT-4V [22].
 From the source image, we extract visual features that can be the object of preservation.
@@
 Finally, AugCLIP is computed by substituting $\mathbf{v}$ to Eq. 3.
-Table 1: Benchmark Datasets and Editing Models.
-Benchmark dataset type that corresponds to different editing scenarios.
-For each dataset, we use different types of editing models that are tailored for each scenario.
-$\dagger$ Kwon et al. (2022); Kim et al. (2022); Kim & Ye (2021); Patashnik et al. (2021)
-$\ddagger$ Hertz et al. (2023); Couairon et al. (2022); Brooks et al. (2022); Hertz et al. (2022); Kawar et al. (2022)
-$\ast$ Li et al. (2024a); Kumari et al. (2023); Wei et al. (2023)
+Table 1: Benchmark Datasets and Editing Models Used for Evaluating the Performance of Metrics.
 CelebA EditVal TEdBench MagicBrush DreamBooth
 Dataset Types Facial Attribute General Object Object Centric Local Region Editing Personalized Generation
-Editing Models $\dagger$ $\ddagger$ $\ddagger$ $\ddagger$ $\ast$
+Editing Models [17, 14, 13, 23] [9, 5, 2, 8, 12] [9, 5, 2, 8, 12] [9, 5, 2, 8, 12] [18, 16, 27]
 Dataset Size 50 648 100 1053 50
@@
 For our experiments, we employ a pre-trained CLIP-ViT 16/B model for CLIP-based metrics.
-Source and target attributes are generated using GPT-4V (OpenAI, 2023).
+Source and target attributes are generated using GPT-4V [22].
 Further details on prompting the source and target attributes are in the Appendix.
@@
 Benchmark Datasets and Editing Models.
-We evaluate AugCLIP and existing metrics across several text-guided image editing benchmarks, including TEdBench (Kawar et al., 2022), EditVal (Basu et al., 2023), MagicBrush (Zhang et al., 2024), DreamBooth (Ruiz et al., 2023), and CelebA Liu et al.
-(2015).
+We evaluate AugCLIP and existing metrics across several text-guided image editing benchmarks, including TEdBench [12], EditVal [1], MagicBrush [28], DreamBooth [26], and CelebA [20].
 Each benchmark dataset represents varying editing scenarios as in Tab.
 1.
-For each benchmark dataset, we employ multiple editing models (Kwon et al., 2022; Kim et al., 2022; Kim & Ye, 2021; Patashnik et al., 2021; Hertz et al., 2023; Couairon et al., 2022; Brooks et al., 2022; Hertz et al., 2022; Li et al., 2024a; Kumari et al., 2023; Wei et al., 2023) as specified in Tab.
-1.
-For CelebA, which require facial attribute editing, we employ domain-specific editing models pre-trained on facial dataset.
-For EditVal, TEdBench, and MagicBrush, we employ generic-purpose editing models.
-For DreamBooth, we use presonalized generation models.
-More details on benchmark datasets are provided in the Appendix.
+For each benchmark dataset, we employ multiple editing models [17, 14, 13, 23, 9, 5, 2, 8, 18, 16, 27] as specified in Tab.
+1 for suitable type of benchmark dataset.
+Details on benchmark datasets are provided in the Appendix.
 5.1 Correlation with Human Judgment
@@
 A high level of alignment between preservation-centric (L2, LPIPS, DINO, SC, CLIP-I) and human judgment is expected because these benchmarks do not significantly deviate from the source image.
+Moreover, per-model evaluations are presented alongside human evaluations in Tab.
+4.
+Unlike other metrics, AugCLIP perfectly aligns with human rankings on each dataset.
 Table 2: User Study Statistics.
@@
 AugCLIP 0.883 0.831 0.857
+Table 4: Rank of various text-guided editing models. Highlighted in green if a metric aligns with human evaluations; otherwise, highlighted in red.
+Dataset Models Rank $\mathrm{CLIP}_{\text{dir}}$uparrow LPIPS $\downarrow$ AugCLIP $\uparrow$ Human $\uparrow$
+DreamBooth ELITE 1 0.1132 2 71.38 2 0.7642 1 0.8478
+BlipDiffusion 2 0.0836 3 70.88 1 0.7579 2 0.6525
+CustomDiffusion 3 0.1348 1 73.84 3 0.6156 3 0.0263
+EditVal P2P 1 0.1771 3 15.04 1 0.8521 1 0.6133
+InstructPix2Pix 2 0.1774 2 25.75 3 0.8242 2 0.4855
+DiffEdit 3 0.2272 1 20.41 2 0.8155 3 0.3214
+CelebA StyleCLIP 1 0.0376 2 27.04 1 0.8484 1 0.6831
+Multi2One 2 0.0414 1 27.95 2 0.8152 2 0.5469
+Asyrp 3 -0.0001 3 36.98 3 0.7750 3 0.3197
 5.2 Ground Truth Editing Selection Test
@@
 In Tab.
-4, AugCLIP is on par with or even better than baseline metrics over all datasets.
+5, AugCLIP is on par with or even better than baseline metrics over all datasets.
 This emphasizes the balanced evaluation criteria of AugCLIP, which previous metrics fail to abide by.
@@
 Table 6: Human Alignment Score ${\bm{s}}_{\text{2AFC}}$ on Various Text-guided Image Editing Scenarios.
-Pos. Add Obj. repl. Alter Parts Background Texture Color Action Style
+Pos. Add Obj. repl. Alter Parts Background
 […]
-AugCLIP 1.0 0.75 0.838 1.0 0.742 1.0 1.0 0.647
+AugCLIP 1.0 0.75 0.838 1.0
+Texture Color Action Style
+[…]
+AugCLIP 0.742 1.0 1.0 0.647
+Design of Weighting
+To support the rationale behind weighting, we design an experiment [^2] Two attributes are compared in terms of relative importance in the same context.
+A total of 10 participants responded in 2AFC format on 40 editing scenarios from the DreamBooth dataset.
+that measures pearson correlation with human evaluation on attribute importance, reported in Tab.
+7.
+Compared to the naïve weighting strategy mimicking $\mathrm{CLIP}_{\text{dir}}$ by measuring cosine similarity between the attribute and the text direction $\Delta T$ (Eq.
+1), AugCLIP weighting (Eq.
+6) demonstrates impressive alignment with human evaluations.
+Table 7: Pearson correlation between weighting strategies and human evaluations on attribute importance.
+Attribute Importance Source Target
+Naïve Weighting w/ $\Delta T$ 0.394 0.637
+AugCLIP Weighting (Eq. 6) 0.794 0.883
 […]
@@
 In Tab.
-8, we compare short (under 5 words) and long (over 5 words), as well as cases with a fixed number (10, 20, or 30) of attributes.
+10, we compare short (under 5 words) and long (over 5 words), as well as cases with a fixed number (10, 20, or 30) of attributes.
 We observe that using short attributes tends to outperform using long ones, likely because they focus on a single visual aspect, avoiding the potential entanglement of multiple aspects in a single attribute.
@@
 short - 0.883 0.831 0.857 0.570 0.889
-Description Randomness
+Equivariance of AugCLIP under Description Randomness
 Using GPT-4V for description extraction introduces slight randomness across different seeds.
-Tests with five seeds show low variance ($0.0197$) in human alignment score (${\bm{s}}_{\text{2AFC}}$), indicating that the description generation process does not affect the scoring.
+To demonstrate near-equivalence quantitatively, we designed a randomness test on the attribute extraction process using the TEdBench dataset, varying five different seeds (examples in Appendix A.
+2).
+ROUGE-L score among attributes extracted by different random seeds is 0.7590 (1 indicates maximum similarity), and the variance of AugCLIP in terms of human alignment is negligible, with a value of 0.0197.
 We also demonstrate in the Appendix that extracted attributes with different seeds do not vary.
@@
 This process takes around 12.3 seconds for attribute generation and 0.15 seconds for score computation.
+This includes the non-reducible CLIP loading time (3.61 sec), which is required by other CLIP-based metrics, and the maximum attribute extraction time (8.62 sec).
 We will provide extracted attributes for all benchmark datasets to reduce this overhead, facilitating the evaluation process for future usage.
@@
 Experiments on five benchmark datasets show that AugCLIP outperforms existing metrics, especially in challenging tasks, offering a more reliable tool for evaluating edits while preserving core attributes.
+Acknowledgements
+This work was supported by the National Research Foundation of Korea (NRF) grants (RS-2023-00209060 A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning), Institute for Information & communications Technology Planning & Evaluation(IITP) grants (No.2022-0-00713 Meta-learning applicable to real-world problems, RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)) funded by the Korea government(MSIT).
 Appendix A Discussions
@@
 In Tab.
-9, we test three variants of optimization objective for deriving the classifier function $g(x)=\mathbf{w}^{T}x+b$.
+11, we test three variants of optimization objective for deriving the classifier function $g(x)=\mathbf{w}^{T}x+b$.
 Since source and target descriptions encoded into CLIP are separable by a simple linear function, we set the hyperplane as $\mathbf{w}^{T}x+b=0$.
@@
 We have emonstrate in Tab.
-9, that this SVM objective is the best option.
+11, that this SVM objective is the best option.
 Moreover, the reasons for choosing this objective are threefold.
@@
 The average percentage of source and target attributes that are wrongly classified by the hyperplane across five benchmark datasets is reported in Tab.
-9, denoted as ‘Average Misc.’.
+11, denoted as ‘Average Misc.’.
 We observe that simple linear decision function $g(x)$ shows a small misclassification rate, $1.35\%$, which signifies its ability to separate the source and target distributions.
@@
 UMAP Visualization.
-Additionally, we visualize that CLIP features of source and target attributes can be separated by a linear hyperplane in 2D projected space using UMAP (McInnes et al., 2018) in Fig.
+Additionally, we visualize that CLIP features of source and target attributes can be separated by a linear hyperplane in 2D projected space using UMAP [21] in Fig.
 8, in which randomly chosen subset of TEdBench samples are plotted.
@@
 Table 12: Assets Employed in Our Experiments. List of pre-trained models, benchmark datasets, and metrics employed in this paper.
-Category Asset URL
-Benchmarks CelebA (Liu et al., 2015) https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
-TedBench (Kawar et al., 2022) https://github.com/imagic-editing/imagic-editing.github.io/tree/main/tedbench
-EditVal (Basu et al., 2023) https://github.com/deep-ml-research/editval_code
-DreamBooth (Ruiz et al., 2023) https://github.com/google/dreambooth
-MagicBrush (Zhang et al., 2024) https://github.com/OSU-NLP-Group/MagicBrush
-Editing Models TEdBench EditVal MagicBrush InstructPix2Pix (Brooks et al., 2022) https://github.com/timothybrooks/instruct-pix2pix
-DiffEdit (Couairon et al., 2022) https://github.com/Xiang-cd/DiffEdit-stable-diffusion.git
-Prompt-to-Prompt (Hertz et al., 2022) https://github.com/google/prompt-to-prompt.git
-DDS (Hertz et al., 2023) https://github.com/google/prompt-to-prompt/blob/main/DDS_zeroshot.ipynb
-Imagic (Kawar et al., 2022) https://github.com/huggingface/diffusers/tree/main/examples/community#imagic-stable-diffusion
-CelebA DiffusionCLIP (Kim & Ye, 2021) https://github.com/gwang-kim/DiffusionCLIP
-Multi2One (Kim et al., 2022) https://github.com/akatigre/multi2one
-Asyrp (Kwon et al., 2022) https://github.com/kwonminki/Asyrp_official
-StyleCLIP (Patashnik et al., 2021) https://github.com/orpatashnik/StyleCLIP
-DreamBooth Custom Diffusion  (Kumari et al., 2023) https://github.com/adobe-research/custom-diffusion
-BLIP-Diffusion (Li et al., 2024a) https://github.com/salesforce/LAVIS/tree/main/projects/blip-diffusion
-ELITE (Wei et al., 2023) https://github.com/csyxwei/ELITE
-Metrics CLIP (Radford et al., 2021) https://github.com/openai/CLIP
-LPIPS (Zhang et al., 2018) https://pypi.org/project/lpips/
-SC (Kim & Ye, 2021) Implemented by the Authors
 C.2 Source Description Generation
@@
 In Tab.
-6, we demonstrate that the simple strategy of directly augmenting $\mathrm{CLIP}_{\text{dir}}$ with the source and target attributes fails to outperform AugCLIP.
+8, we demonstrate that the simple strategy of directly augmenting $\mathrm{CLIP}_{\text{dir}}$ with the source and target attributes fails to outperform AugCLIP.
 Additionally, we show the effect of applying the weighting strategy () of Eq.
 6 when aggregating attributes into $\mathrm{CLIP}_{\text{dir}}$ in Tab.
-11.
+13.
 Note that augmentation of $\mathrm{CLIP}_{\text{dir}}$ without weighting () is already reported in Tab.
-6.
+8.
 Formally, for augmenting the source text $T_{\text{src}}$ with source attributes in $\mathcal{D}_{S}$, directional CLIP similarity is redefined as
@@
 AugCLIP 0.883 0.831 0.857 0.570 0.889
-Recently, GPT-4V (OpenAI, 2023) has been employed in evaluating various tasks, including text-guided image editing, text-to-image generation, and image quality assessment.
+Recently, GPT-4V [22] has been employed in evaluating various tasks, including text-guided image editing, text-to-image generation, and image quality assessment.
 Since GPT-4V is one of the best-performing multi-modal large language models, we test the ability of GPT-4V’s effectiveness in evaluating the quality of text-guided edited images.
@@
 As shown in Tab.
-12, GPT-4V outperforms AugCLIP in tasks such as EditVal and TEdBench, which involve simple edits like modifying a single object’s attribute.
+14, GPT-4V outperforms AugCLIP in tasks such as EditVal and TEdBench, which involve simple edits like modifying a single object’s attribute.
 In contrast, our proposed metric, AugCLIP, effectively captures minor differences by augmenting attributes of the source image and target text and shows better performance in other benchmarks with complex scenarios.
-This finding is consistent with prior research (Zhang et al., 2023), which suggests that GPT-4V struggles to differentiate between images with subtle differences.
+This finding is consistent with prior research [30], which suggests that GPT-4V struggles to differentiate between images with subtle differences.
 D.4 Additional Examples on Problem 1 of Directional CLIP Similarity
@@
 […]
-To assess directional CLIP similarity’s capability to focus on the image regions modified following the target text rather than peripheral or unchanged regions, we use the relevancy map (Chefer et al., 2021), ${\bm{R}}$.
+To assess directional CLIP similarity’s capability to focus on the image regions modified following the target text rather than peripheral or unchanged regions, we use the relevancy map [4], ${\bm{R}}$.
 The relevancy map visualizes the transformer’s attention on an image corresponding to a given text depending on their cosine similarity.
@@
 Unlike $\mathrm{CLIP}_{\text{dir}}$, AugCLIP measures the cosine similarity between the estimated well-edited image and the edited image, rather than between an image and text.
-As a result, the relevancy map, which requires a direct comparison of the image and text, cannot be applied to AugCLIP.
+As a result, the relevancy map, which requires direct comparison of the image and text, cannot be applied to AugCLIP.
 Appendix E Qualitative Results
@@
 Additionally, we provide qualitative samples from the Ground Truth Selection Test, reported in Tab.
-4, using the TEdBench and MagicBrush datasets (Fig.
+5, using the TEdBench and MagicBrush datasets (Fig.
 18, 19 and Fig.